OpenStudy (anonymous):

You use an inclined plane to move furniture into a truck. You perform 105 kJ of work, but only do 88 kJ of useful work. What is the efficiency of the inclined plane?

OpenStudy (anonymous):

\[\frac{88~\text{kJ}}{105~\text{kJ}}\]

OpenStudy (anonymous):

what would it be out of 83.8% 34.6% 12.3% 119.0% @broken_symmetry

OpenStudy (anonymous):

\[83.8\% = \frac{83.8}{100} \approx \frac{88}{105}\] You're dividing by a number very close to 100, so the numerator is going to be roughly the same in percentage.

OpenStudy (anonymous):

thanks so much @broken_symmetry